OpenStudy (anonymous):

Use g=10 N/kg in this assignment. 1. What is the gravitational energy (relative to the unstretched surface of the trampoline) of the 20 kg ball at its apex 2 m above the trampoline? 2. What is the kinetic energy of the ball just before impacting the trampoline? 3. At maximum stretch at the bottom of the motion, what is the sum of the elastic and gravitational energy of the ball? 4. What conclusions can you draw from the answers found above? for #2 would the answer be 0 J? and I also need help with #4

OpenStudy (anonymous):

for number 2, it is asking for the KE before the ball hits the trampoline, so it should not be equal to zero J.

OpenStudy (anonymous):

so when the ball is at the top : gravitational potential energy=mgh kinetic energy=0J elastic potential energy=0J Eg + 0 + 0=total energy JUST Before it hits the trampoline: 0 + Ek + 0 =total energy when the ball hits the trampoline at a maximum stretch Eg + 0 + Ee = total energy